The general solution of the equation is y'+(cotx)y=xcscx
1
Expert's answer
2019-10-14T12:00:12-0400

"y'+(cotx)y=xcscx
\newline
\frac{dy}{dx}+(cotx)y=(cscx)x"


This equation is a linear equation of the form


"\frac{dy}{dx}+P(x)y=Q(x)"


To solve this equation, we multiply both sides by the integrating factor I(x):


"I(x)=e^{\int P(x)dx}=e^{\int cotxdx}
\newline
\int cotxdx = \int \frac{cosx}{sinx}dx = \int \frac{d(sinx)}{sinx} = ln|sinx|+c"


We take c=0. Then we get:


"I(x)=e^{ln|sinx|}=|sinx|"


Let's multiply both sides by sinx (absolute value doesn't really change anything):


"sin(x)\frac{dy}{dx}+sinx\cdot cotx \cdot y=x(cscx \cdot sinx)
\newline
sinx\frac{dy}{dx}+cosx \cdot y=x
\newline
\frac{d}{dx}(y\cdot sinx)=x"


Integrating both sides we get the answer:


"y\cdot sinx=\frac{x^2}{2}+c
\newline
y=\frac{\frac{x^2}{2}+c}{sinx}"